Explain ohm s law?

1
Expert's answer
2021-12-30T11:59:34-0500

Ohm law :-

the current through a conductor between two points is directly proportional to the voltage drop across the two points.



Voltage perportional to flow ofcurrent in wire

"V=IR"

R=perportional constant

Where

R=resistance

V=voltage

I=current

Ohm's for graph



Slop in tringle

"tan\theta=\frac{∆y}{∆x}"

"tan\theta=\frac{∆I}{∆V}"

"m=tan\theta=\frac{∆I}{∆V}"

"m=\frac{1}{R}"

"\frac{1}{R}=\frac{∆I}{∆V}"

Now We can written as

∆V=R∆I

This is ohm's law